OpenStudy (anonymous):

use the intercepts to graph the equation x-6=y

OpenStudy (anonymous):

y=x-6 slope=+1 try fixing different values for x and y like say if x is 2 y is -4 then plot

OpenStudy (anonymous):

y=-4

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

slope = 1 x-6=y set x=0 find y intercept 0-6=y .............(0,-6) set y= 0 find x intercept x-6=0 x = 6 (6, -6) x intercept is: 6 y intercept is: -6 y=-4 , x= 6 is wrong answer http://www.mathway.com/answer.aspx?p=calg?p=x-6SMB01y?p=105?p=?p=?p=?p=?p=0?p=?p=0?p=?p=